The amplitude of an oscillator decreases to 55.4 % of its initial value in 14.0 s .
Answer #1

Let the time constant be 'tau' then we have

A(t) = A(0)*e^-(t/tau).

where where t is the time in seconds, A(t) is the amplitude at time t, A(0) is the original amplitude, Using the given data,

It is given that.
0.554 = e^(-14/tau) or
ln(0.554) = -14/tau or

tau = -14/ln(0.554) = 23.73 s

tau = 23.73 s